OpenStudy (anonymous):

Estimate the heat of neutralization of H3PO4 in kJ/mol given these data: Heat transferred = 2.2 x 102 J Moles of H3PO4 = 1.5 x 10−3 mol Moles of NaOH = 5.0 x 10−3 mol

OpenStudy (anonymous):

1.) write out the balance equation 3NaOh+H3PO4->Na3PO4+3H2O 2.) You are given everything needed to calculate q=heat transfer=2.2*10^2, H3PO4 moles= 1.5*10^-3, NaOH moles=5.0*10^-3 3.) equation is deltaHneutraliztion=q/Moles of limiting reagent H3PO4 is limiting reagent because lowest moles, and is used up first 4.) Now plug in variables DeltaH=2.2*10^2(1.5*10^3)= 146.67kj/mole Notice we had to convert J to kj,